Selected transactions of Shadrach Computer Corporation during November and December of 2019 are as follows: November TransactDecember Transactions: Dec. Twenty proofs of purchase were returned from the November 12 sale. 14 29 An individual filed suitChart of Accounts ASSETS REVENUE 111 Cash 411 Sales 121 Accounts Receivable 132 Inventory of Premiums EXPENSES 152 Prepaid InEQUITY 311 Common Stock 331 Retained EarningsPrepare journal entries to record the November transactions of Shadrach Computer Corporation for 2019. Include year-end accruPAGE 11 GENERAL JOURNAL DATE ACCOUNT TITLE POST. REF. DEBIT CREDITPrepare journal entries to record the December transactions of Shadrach Computer Corporation for 2019. Include year-end accruGENERAL JOURNAL DATE ACCOUNT TITLE POST. REF. DEBIT CREDIT
Answer #1
Date Accounts title Ref Dr Cr
01-Nov Cash 111           51,975
Discount On Note payable 261             2,025
Note Payable 260           54,000
31-Dec Interest expense 2 months 540             1,350
Discount On Note payable 261             1,350
09-Nov Accounts Receivable 121        6,24,000
Sales Revenue 411        6,24,000
Warranty expense 551           13,800
Estimated Warranty Liability 240           13,800
12-Nov Accounts Receivable 120*290 121           34,800
Sales Revenue 411           34,800
Premium expense 552             1,440
Estimated Premium Liability 241             1,440
20-Nov Estimated Warranty Liability 240             2,500
Cash 111             2,500
30-Nov Salaries Expenses (190*12*100*(1/12)) 521           19,000
Liability for Compensated Absences 222           19,000
30-Nov Salaries Expenses 521        4,40,000
Federal income taxes witholding payable
(0.20 x $440,000)
225           88,000
F.I.C.A taxes payable
(0.08 x $440,000)
224           35,200
Cash (Balancing Figure) 111        3,16,800
30-Nov Payroll Tax Expense 522           35,200
F.I.C.A taxes payable
(0.08 x $440,000)
224           35,200
14-Dec Estimated Premium Liability 241                 300
Cash 111                 300
29-Dec Loss from accident 561             1,400
Estimated Liability from Lawsuit 245             1,400
31-Dec Salaries Expenses (190*12*100*(1/12)) 521           19,000
Liability for Compensated Absences 222           19,000
31-Dec Salaries expense 521           13,200
Bonus payable 221           13,200
(560000*.7= (392000-260000)*10%=
31-Dec Salaries Expenses 521        4,33,000
Federal income taxes witholding payable
(0.20 x $433,000)
225           86,600
F.I.C.A taxes payable
(0.08 x $433,000)
224           34,640
Cash (Balancing Figure) 111        3,11,760
31-Dec Payroll Tax Expense 522           34,640
F.I.C.A taxes payable
(0.08 x $440,000)
224           34,640
31-Dec Salaries' Expenses 521           56,000
Bonus Payable 221           56,000